# I INTRODUCTION

### v1

@@ I INTRODUCTION
The evolution towards 6G is expected to enable a new generation of applications, for example in the area of cyber-physical systems [Karapantelakis2024, liu2017review], demanding ultra-reliable, real-time, and low-latency communication. For such applications, it is crucial to assess performance indicators such as One-Way Delay (OWD) [owd] and Round Trip Time (RTT) [rtt], as even minor disturbances in the communication between two network functions could lead to severe operational and safety risks. Unfortunately, there is complexity associated with assessing the performance outcome of a User Equipment (UE) as it is highly influenced by and dependent on multiple factors such as network configuration, signal conditions, competing traffic, exogenous processes in the UE, and variations in UE hardware [ansari2022performance]. A promising approach, explored in both academia and industry, is based upon Machine Learning (ML) where the performance outcome of a UE, such as the OWD [riihijarvi2018machine, mostafavi2023data, rao2022prediction], is predicted based upon statistics available to the network operator. Accurate performance prediction, not only with respect to point estimates but also with respect to distributions and tails [mostafavi2023data, flinta2020predicting, samani2021conditional], is essential for proactive service assurance, troubleshooting, verification, and performance optimization.
@@
Previous research has demonstrated the feasibility of modeling delay; however, it has also shown that model performance deteriorates over time due to distributional shifts in data. These shifts are due to factors such as previously unseen user behavior patterns (e.g., traffic fluctuations, mobility) and variations in UE chip sets [rao2024generalizable]. Various mitigation strategies have been explored, including transfer learning and domain adaptation [larsson2023domain], and semi-supervised learning [rao2025self]. More broadly, it is well established that network management models degrade over time due to the dynamic nature of networks and their underlying compute resources [moradi2019performance]. The challenge of maintaining network management models over time has also been identified in standardization, and thus there is an increased interest in incorporating automated ML workflows and continuous operations into the network architecture.
Despite recent advancements in model generalization, a significant challenge remains: sustaining model performance over extended periods. This challenge is illustrated in Fig. 1, where a network environment evolves over time, introducing distributional shifts that can degrade model performance. Each time step in this dynamic environment can be considered as a continual learning (CL) task, new data is observed and thereafter integrated with the existing generalized model through CL, producing an updated generalized model. The key challenge lies in balancing model plasticity with learning stability [dohare2024loss] - ensuring that new behaviors in the network environment are learned while preserving previously acquired knowledge. In other words, preventing catastrophic forgetting in ML [McCloskey1989CatastrophicII]. This paper specifically addresses this challenge for OWD prediction models, as discussed above, and proposes a novel approach based on CL. The method leverages generative replay [Shin2017ContinualLW] and extends such approaches with a multi-generator framework, where generators are designed and maintained using domain knowledge of the network and its UEs. Generative replay, specifically the proposed multi-generator approach, not only mitigates catastrophic forgetting, it also improves performance on the distribution tail as well as enabling network operators to reduce the storage requirements for previously seen data samples.
@@ II-A BACKGROUND
The objective of this paper is to establish a robust approach for training and maintaining a ML model that predicts OWD as experienced by a UE. This prediction is based on measurements taken with high frequency from the baseband in the Radio Access Network (RAN). Although our previous work [rao2024generalizable] addressed some aspects of OWD model generalization, the expanded aim of this paper is to prevent catastrophic forgetting, and thus ensure robustness during long-term maintenance of OWD models.
@@ II-C PROBLEM STATEMENT
The deep generative replay framework [Shin2017ContinualLW] presents a promising approach to CL, addressing the limitations associated with the data replay. While potentially effective, we argue that the generative replay itself can suffer from catastrophic forgetting. As the number of tasks increases over time, with the growing diversity resulting from introduction of new tasks, the generative model tends to capture merely the modalities in the data that represent the bulk of the underlying data distribution and fail to capture all the modalities. We refer to this phenomenon as mode collapse that is when a modality that was present once collapses into a more dominant modality.
@@ III-A1 Deep generative replay
The model architecture of the deep generative replay consists of a generative model, generator, and a task solving model, solver. Shin et al. in [Shin2017ContinualLW] refer to this dual model architecture as scholar, defined by $\mathcal{H}=\{G,S\}$} where $G$ denotes the generator and $S$ denotes the solver. The generator is a parameterized generative model, with a parameter set denoted by $\phi$, that is learned to produce real-like samples of the input features, and the solver is a parameterized predictive model of the task with a parameter set denoted by $\theta$.
@@ III-A2 Tabular Variational Auto-Encoders
TVAE is the state-of-the-art variational auto-encoder (VAE) for tabular data generation [Xu2019ModelingTD]. It is specifically designed for tabular data, which is typically composed of a mix of both discrete and continuous features. Continuous features may have multi-modal and non-Gaussian distributions whereas discrete features are sometimes imbalanced making the modeling difficult.
The basic scheme of a VAE composes of an encoder and a decoder [Kingma2013AutoEncodingVB]. The encoder compresses the input $\boldsymbol{x}$} into the latent space. The decoder receives as input the information sampled from the latent space and produces $\boldsymbol{x}^{\prime}$} as similar as possible to $\boldsymbol{x}$}. In TVAE, the modeling for encoder is similar to conventional VAE. The decoder is designed specially so that the probability distribution of data can be modeled accurately; refer to [Xu2019ModelingTD] for details. Also, a mode-specific normalization is designed to deal with features with complicated distributions. This architecture allows TVAE to generate data that closely match the distribution of real tabular data, making it more suitable for applications like data imputation, synthetic data generation, and tabular data modeling.
@@ III-B1 Generator relevance determination
Let $\boldsymbol{a}=(a_{1},\ldots,a_{J})^{\top}$}, for all $a_{j}\in\{0,1\}$}, denote a binary vector of $J$ data configurations at a given task condition where each element specifies a different configuration, referred to as the task configuration vector. In [rao2024generalizable], Rao et al. showed that OWD distribution is significantly dependent upon the UE type. Therefore, in our use case of OWD prediction, the elements of the task configuration vector represent UE device type; as an example, for three different UE types as described in Table I, the task configuration vector is expressed by a 3-dimensional vector with binary elements with the three elements corresponding to the UE types,
@@ IV A 5G TESTBED FOR CREATION OF ONE-WAY DELAY DATASETS
To evaluate the multi-generator generative replay approach for robust OWD prediction in future 6G networks, we leverage an in-house 5G-mmWave testbed [rao2024generalizable] to create datasets consisting of base station features and ground-truth OWD values. In the following, while we briefly describe the testbed, we focus on the experiments conducted in support of evaluating our approach.
The testbed is a 5G NSA system build upon commercially available Ericsson StreetMacro 6701 device that implements 5G NR, and an Ericsson Radio Dot 2243 device that implements 4G LTE. The StreetMacro does analog beamforming using one horizontal and one vertical beam (1 × 200 MHz dual polarized beams). We set the 4G LTE eNB to operate on band B3 (1800 MHz frequency, with 5 MHz bandwidth), and 5G NR gNodeB (gNB) on band n257 (28 GHz frequency, with 100 MHz bandwidth). The testbed resides in an indoor testbed area. The floor plan and associated positions of the 5G NR and 4G LTE radios, is illustrated in Fig. 3. For details regarding the setup we refer to [rao2024generalizable].
Similarly to [rao2024generalizable], we measure OWD, between the UE and a receiver in the network, with three different types of 5G UE devices, with different manufacturers and chipsets, referred to as UE1, UE2, UE3. The OWD measurements are conducted using TWAMP [hedayat2008two] probe packets of sizes 100 and 1400 bytes. An additional UE is configured for creation of network traffic scenarios, where that UE is introducing competing uplink traffic into the network, ranging from 0 to 60 Mbps, in steps of 10 Mbps. A UE is either stationary (position 1, 2, or 3), or moving according to a set of predefined movement patterns (rectangular, zigzag, and linear). The experiment parameters are summarized in Table I.
@@
In addition to collection of ground-truth OWD, the testbed was configured with multiple measurement points to extract network metrics [rao2024generalizable]. More specifically, we extract logs from the gNB corresponding to beam forming, UE connectivity, and uplink (UL)/ downlink (DL) scheduling events [3gpp.38.214, 3gpp.38.215, 3gpp.38.321], on the sub-millisecond scale. Of a much larger set available, we choose to monitor and use 103 metrics.
@@ V-A DESIGNING CL TASK SEQUENCES
In [rao2024generalizable], Rao et al. showed that various configuration settings such as UE device type and the UE movement pattern can result in distributional shift both at the input space as well as in the task space (OWD). Inspired by this, we create a series of CL tasks by utilizing configuration settings from various experiments to define the tasks. In particular, each task is constructed based on the UE type, along with its position and movement pattern as defined and illustrated in Table I and Fig. 3. Accordingly, two groups of CL task sequences are constructed which are shown in Table II.
@@ 8 UE 1,2,3 UE 1 3 2 1 3 2 1 3 2 1 3 2 1 3 2 1 3 2
Generator TVAE The TVAE Synthesizer in python library SDV [SDV] is used to train a TVAE model and generate synthetic data.
@@ V-B1 Baseline approaches
As in [Maltoni2018ContinuousLI], Naïve method could be used as the lower bound baseline for stability evaluation, in which the model is trained continuously with data of each task, without any particular framework to control forgetting. That is, the solver is finetuned using data from the new task, without access to the data from the previous tasks.
@@
Cumulative method is used to evaluate the upper bound of forgetting in CL as in [Maltoni2018ContinuousLI]. For each task, it accumulates all data from previous tasks and data of current task, re-train the solver from scratch. There is no forgetting as cumulative method could access and use all the previous data. However, it is often infeasible in real world applications, due to constrains in data storage and privacy.
@@
It is illustrated in the state-of-the-art deep generative replay algorithm[Shin2017ContinualLW] that a VAE could be used as the generator. We evaluate its performance on our cases. The architecture of the VAE is shown in Table III. It is worth noting that we experimented with different numbers of layers and layer sizes where all of which produced similar results.
@@ V-C PERFORMANCE EVALUATION
As the prediction of OWD is a regression task, the performance of the model is evaluated using a standard percentage error metric known as mean absolute percentage error (MAPE), aligning with our prior study [rao2024generalizable], thereby facilitating consistent benchmarking and comparative analysis. MAPE is defined as:
@@
To evaluate the performance of the selected CL algorithms, we split the dataset for each CL task into training and test sets (70% for training). Training is done sequentially with the training set of each task according to the task sequence. Performance is then evaluated using the test sets for all $I$ tasks. After the model finishes learning of all $I$ tasks, we get the result matrix $R\in\mathbb{R}^{I,I}$}, where $R_{i,j}$ is the performance of the model on test set of task $T_{j}$ after training with the training set of task $T_{i}$. Specifically for our evaluations, the model performance $R_{i,j}$ is MAPE values. Inspired by the two CL evaluation metrics Average Accuracy (ACC) and Backward Transfer (BWT) in [LopezPaz2017GradientEM], we define metrics Average MAPE (AveMAPE) and Forgetting(F) for our regression task OWD prediction, where
@@ VIII RELATED WORK
CL methods have been proposed to improve various aspects of ML and has been categorized in different ways [Wang2023ACS][lesort2019continuallearningroboticsdefinition][DeLange2019ACL]. In [Wang2023ACS], based on how task specific information is stored and used throughout the sequential learning process, the authors distinguish five major categories, which are Regularization-Based approach, Replay-Based approach, Optimization-Based approach, Representation-Based approach and Architecture-Based approach. The different categories come with different properties related to energy usage as described in [trinci2024green], where the authors compare different CL approaches in terms of energy consumption. Here the Representation-Based approaches score highest on energy efficiency but other approaches such as Replay-Based still outperform the naive baseline of joint training.
@@
The focus of our paper is generative replay or pseudo-rehearsal, which is a sub-direction in the Replay-Based approach. Instead of storing old training samples, generative replay requires training a generative model to replay generated data. Compared with other approaches, one benefit of generative replay is that the generative model makes it possible to provide data samples from previous tasks for future needs, with much less memory and feasible for privacy concerns. As it is presented in Section III, DGR[Shin2017ContinualLW] establishes a framework that the learning of each new task is coupled with replaying generated data sampled from one generative model, ensuring the retention of previously acquired knowledge. MeRGAN [Wu2018MemoryRG] incorporates a memory replay mechanism to prevent catastrophic forgetting, which is further enforces through either joint training or replay alignment.
@@
CL approaches in various categories have advantages and work best within specific scenarios. Between the categories, approaches are often orthogonal with respect to each other. Hence, generative replay could be hybrid with other CL strategies. To mitigate catastrophic forgetting of generative models, VCL[Nguyen2017VariationalCL] uses a weight regularization based variational Bayesian approach, and maintaining a distribution over parameters and updating it using evidence lower bound (ELBO). DGM [Ostapenko2019LearningTR] is inspired by biological synaptic plasticity principles, incorporating mechanisms like memory consolidation and task-specific adaptations.
@@
For pseudo-rehearsal, the generative models could be of various types, such as generative adversarial networks (GANs) and variational autoencoder (VAE). GANs are typically used in scenarios where high-quality, realistic samples are required, such as image generation, super-resolution, and artificial data generation [Ostapenko2019LearningTR]. GANs are highly flexible and can be adapted to various types of generative tasks, but can be harder to control due to the adversarial nature. VAEs are more suited for tasks where a structured latent space and stable training are important [Kemker2017FearNetBM] [Riemer2017ScalableRF]. Since VAEs have an explicit, interpretable latent space, they are often more useful in applications where understanding the latent factors of data is important. Besides GANs and VAE, the novel approach DDGR [Gao2023DDGRCL] adopts a diffusion model as the generator and calculates an instruction-operator through the classifier to instruct the generation of samples.

### v2

@@ I INTRODUCTION
The evolution towards 6G is expected to enable a new generation of applications, for example in the area of cyber-physical systems [11, 17], demanding ultra-reliable, real-time, and low-latency communication. For such applications, it is crucial to assess performance indicators such as One-Way Delay (OWD) [4] and Round Trip Time (RTT) [10], as even minor disturbances in the communication between two network functions could lead to severe operational and safety risks. Unfortunately, there is complexity associated with assessing the performance outcome of a User Equipment (UE) as it is highly influenced by and dependent on multiple factors such as network configuration, signal conditions, competing traffic, exogenous processes in the UE, and variations in UE hardware [5]. A promising approach, explored in both academia and industry, is based upon Machine Learning (ML) where the performance outcome of a UE, such as the OWD [30, 22, 28], is predicted based upon statistics available to the network operator. Accurate performance prediction, not only with respect to point estimates but also with respect to distributions and tails [22, 7, 31], is essential for proactive service assurance, troubleshooting, verification, and performance optimization.
@@
Previous research has demonstrated the feasibility of modeling delay; however, it has also shown that model performance deteriorates over time due to distributional shifts in data. These shifts are due to factors such as previously unseen user behavior patterns (e.g., traffic fluctuations, mobility) and variations in UE chip sets [27]. Various mitigation strategies have been explored, including transfer learning and domain adaptation [15], and semi-supervised learning [26]. More broadly, it is well established that network management models degrade over time due to the dynamic nature of networks and their underlying compute resources [21]. The challenge of maintaining network management models over time has also been identified in standardization, and thus there is an increased interest in incorporating automated ML workflows and continuous operations into the network architecture.
Despite recent advancements in model generalization, a significant challenge remains: sustaining model performance over extended periods. This challenge is illustrated in Fig. 1, where a network environment evolves over time, introducing distributional shifts that can degrade model performance. Each time step in this dynamic environment can be considered as a continual learning (CL) task, new data is observed and thereafter integrated with the existing generalized model through CL, producing an updated generalized model. The key challenge lies in balancing model plasticity with learning stability [6] - ensuring that new behaviors in the network environment are learned while preserving previously acquired knowledge. In other words, preventing catastrophic forgetting in ML [20]. This paper specifically addresses this challenge for OWD prediction models, as discussed above, and proposes a novel approach based on CL. The method leverages generative replay [32] and extends such approaches with a multi-generator framework, where generators are designed and maintained using domain knowledge of the network and its UEs. Generative replay, specifically the proposed multi-generator approach, not only mitigates catastrophic forgetting, it also improves performance on the distribution tail as well as enabling network operators to reduce the storage requirements for previously seen data samples.
@@ II-A BACKGROUND
The objective of this paper is to establish a robust approach for training and maintaining a ML model that predicts OWD as experienced by a UE. This prediction is based on measurements taken with high frequency from the baseband in the Radio Access Network (RAN). Although our previous work [27] addressed some aspects of OWD model generalization, the expanded aim of this paper is to prevent catastrophic forgetting, and thus ensure robustness during long-term maintenance of OWD models.
@@ II-C PROBLEM STATEMENT
The deep generative replay framework [32] presents a promising approach to CL, addressing the limitations associated with the data replay. While potentially effective, we argue that the generative replay itself can suffer from catastrophic forgetting. As the number of tasks increases over time, with the growing diversity resulting from introduction of new tasks, the generative model tends to capture merely the modalities in the data that represent the bulk of the underlying data distribution and fail to capture all the modalities. We refer to this phenomenon as mode collapse that is when a modality that was present once collapses into a more dominant modality.
@@ III-A1 Deep generative replay
The model architecture of the deep generative replay consists of a generative model, generator, and a task solving model, solver. Shin et al. in [32] refer to this dual model architecture as scholar, defined by $\mathcal{H}=\{G,S\}$} where $G$ denotes the generator and $S$ denotes the solver. The generator is a parameterized generative model, with a parameter set denoted by $\phi$, that is learned to produce real-like samples of the input features, and the solver is a parameterized predictive model of the task with a parameter set denoted by $\theta$.
@@ III-A2 Tabular Variational Auto-Encoders
TVAE is the state-of-the-art variational auto-encoder (VAE) for tabular data generation [36]. It is specifically designed for tabular data, which is typically composed of a mix of both discrete and continuous features. Continuous features may have multi-modal and non-Gaussian distributions whereas discrete features are sometimes imbalanced making the modeling difficult.
The basic scheme of a VAE composes of an encoder and a decoder [13]. The encoder compresses the input $\boldsymbol{x}$} into the latent space. The decoder receives as input the information sampled from the latent space and produces $\boldsymbol{x}^{\prime}$} as similar as possible to $\boldsymbol{x}$}. In TVAE, the modeling for encoder is similar to conventional VAE. The decoder is designed specially so that the probability distribution of data can be modeled accurately; refer to [36] for details. Also, a mode-specific normalization is designed to deal with features with complicated distributions. This architecture allows TVAE to generate data that closely match the distribution of real tabular data, making it more suitable for applications like data imputation, synthetic data generation, and tabular data modeling.
@@ III-B1 Generator relevance determination
Let $\boldsymbol{a}=(a_{1},\ldots,a_{J})^{\top}$}, for all $a_{j}\in\{0,1\}$}, denote a binary vector of $J$ data configurations at a given task condition where each element specifies a different configuration, referred to as the task configuration vector. In [27], Rao et al. showed that OWD distribution is significantly dependent upon the UE type. Therefore, in our use case of OWD prediction, the elements of the task configuration vector represent UE device type; as an example, for three different UE types as described in Table I, the task configuration vector is expressed by a 3-dimensional vector with binary elements with the three elements corresponding to the UE types,
@@ IV A 5G TESTBED FOR CREATION OF ONE-WAY DELAY DATASETS
To evaluate the multi-generator generative replay approach for robust OWD prediction in future 6G networks, we leverage an in-house 5G-mmWave testbed [27] to create datasets consisting of base station features and ground-truth OWD values. In the following, while we briefly describe the testbed, we focus on the experiments conducted in support of evaluating our approach.
The testbed is a 5G NSA system build upon commercially available Ericsson StreetMacro 6701 device that implements 5G NR, and an Ericsson Radio Dot 2243 device that implements 4G LTE. The StreetMacro does analog beamforming using one horizontal and one vertical beam (1 × 200 MHz dual polarized beams). We set the 4G LTE eNB to operate on band B3 (1800 MHz frequency, with 5 MHz bandwidth), and 5G NR gNodeB (gNB) on band n257 (28 GHz frequency, with 100 MHz bandwidth). The testbed resides in an indoor testbed area. The floor plan and associated positions of the 5G NR and 4G LTE radios, is illustrated in Fig. 3. For details regarding the setup we refer to [27].
Similarly to [27], we measure OWD, between the UE and a receiver in the network, with three different types of 5G UE devices, with different manufacturers and chipsets, referred to as UE1, UE2, UE3. The OWD measurements are conducted using TWAMP [9] probe packets of sizes 100 and 1400 bytes. An additional UE is configured for creation of network traffic scenarios, where that UE is introducing competing uplink traffic into the network, ranging from 0 to 60 Mbps, in steps of 10 Mbps. A UE is either stationary (position 1, 2, or 3), or moving according to a set of predefined movement patterns (rectangular, zigzag, and linear). The experiment parameters are summarized in Table I.
@@
In addition to collection of ground-truth OWD, the testbed was configured with multiple measurement points to extract network metrics [27]. More specifically, we extract logs from the gNB corresponding to beam forming, UE connectivity, and uplink (UL)/ downlink (DL) scheduling events [1, 2, 3], on the sub-millisecond scale. Of a much larger set available, we choose to monitor and use 103 metrics.
@@ V-A DESIGNING CL TASK SEQUENCES
In [27], Rao et al. showed that various configuration settings such as UE device type and the UE movement pattern can result in distributional shift both at the input space as well as in the task space (OWD). Inspired by this, we create a series of CL tasks by utilizing configuration settings from various experiments to define the tasks. In particular, each task is constructed based on the UE type, along with its position and movement pattern as defined and illustrated in Table I and Fig. 3. Accordingly, two groups of CL task sequences are constructed which are shown in Table II.
@@ 8 UE 1,2,3 UE 1 3 2 1 3 2 1 3 2 1 3 2 1 3 2 1 3 2
Generator TVAE The TVAE Synthesizer in python library SDV [25] is used to train a TVAE model and generate synthetic data.
@@ V-B1 Baseline approaches
As in [19], Naïve method could be used as the lower bound baseline for stability evaluation, in which the model is trained continuously with data of each task, without any particular framework to control forgetting. That is, the solver is finetuned using data from the new task, without access to the data from the previous tasks.
@@
Cumulative method is used to evaluate the upper bound of forgetting in CL as in [19]. For each task, it accumulates all data from previous tasks and data of current task, re-train the solver from scratch. There is no forgetting as cumulative method could access and use all the previous data. However, it is often infeasible in real world applications, due to constrains in data storage and privacy.
@@
It is illustrated in the state-of-the-art deep generative replay algorithm[32] that a VAE could be used as the generator. We evaluate its performance on our cases. The architecture of the VAE is shown in Table III. It is worth noting that we experimented with different numbers of layers and layer sizes where all of which produced similar results.
@@ V-C PERFORMANCE EVALUATION
As the prediction of OWD is a regression task, the performance of the model is evaluated using a standard percentage error metric known as mean absolute percentage error (MAPE), aligning with our prior study [27], thereby facilitating consistent benchmarking and comparative analysis. MAPE is defined as:
@@
To evaluate the performance of the selected CL algorithms, we split the dataset for each CL task into training and test sets (70% for training). Training is done sequentially with the training set of each task according to the task sequence. Performance is then evaluated using the test sets for all $I$ tasks. After the model finishes learning of all $I$ tasks, we get the result matrix $R\in\mathbb{R}^{I,I}$}, where $R_{i,j}$ is the performance of the model on test set of task $T_{j}$ after training with the training set of task $T_{i}$. Specifically for our evaluations, the model performance $R_{i,j}$ is MAPE values. Inspired by the two CL evaluation metrics Average Accuracy (ACC) and Backward Transfer (BWT) in [18], we define metrics Average MAPE (AveMAPE) and Forgetting(F) for our regression task OWD prediction, where
@@ VIII RELATED WORK
CL methods have been proposed to improve various aspects of ML and has been categorized in different ways [34][16][14]. In [34], based on how task specific information is stored and used throughout the sequential learning process, the authors distinguish five major categories, which are Regularization-Based approach, Replay-Based approach, Optimization-Based approach, Representation-Based approach and Architecture-Based approach. The different categories come with different properties related to energy usage as described in [33], where the authors compare different CL approaches in terms of energy consumption. Here the Representation-Based approaches score highest on energy efficiency but other approaches such as Replay-Based still outperform the naive baseline of joint training.
@@
The focus of our paper is generative replay or pseudo-rehearsal, which is a sub-direction in the Replay-Based approach. Instead of storing old training samples, generative replay requires training a generative model to replay generated data. Compared with other approaches, one benefit of generative replay is that the generative model makes it possible to provide data samples from previous tasks for future needs, with much less memory and feasible for privacy concerns. As it is presented in Section III, DGR[32] establishes a framework that the learning of each new task is coupled with replaying generated data sampled from one generative model, ensuring the retention of previously acquired knowledge. MeRGAN [35] incorporates a memory replay mechanism to prevent catastrophic forgetting, which is further enforces through either joint training or replay alignment.
@@
CL approaches in various categories have advantages and work best within specific scenarios. Between the categories, approaches are often orthogonal with respect to each other. Hence, generative replay could be hybrid with other CL strategies. To mitigate catastrophic forgetting of generative models, VCL[23] uses a weight regularization based variational Bayesian approach, and maintaining a distribution over parameters and updating it using evidence lower bound (ELBO). DGM [24] is inspired by biological synaptic plasticity principles, incorporating mechanisms like memory consolidation and task-specific adaptations.
@@
For pseudo-rehearsal, the generative models could be of various types, such as generative adversarial networks (GANs) and variational autoencoder (VAE). GANs are typically used in scenarios where high-quality, realistic samples are required, such as image generation, super-resolution, and artificial data generation [24]. GANs are highly flexible and can be adapted to various types of generative tasks, but can be harder to control due to the adversarial nature. VAEs are more suited for tasks where a structured latent space and stable training are important [12] [29]. Since VAEs have an explicit, interpretable latent space, they are often more useful in applications where understanding the latent factors of data is important. Besides GANs and VAE, the novel approach DDGR [8] adopts a diffusion model as the generator and calculates an instruction-operator through the classifier to instruct the generation of samples.